5.) A car traveling at 15 m/s passes a van that is starting from rest with an acceleration of 3 m/s². (a) How long will it take the van to catch up with the car? (b) How far has the car traveled when the van catches up with it?
1
Expert's answer
2019-08-27T09:33:46-0400

Let after time "t \space sec" ,van catches up the car.

(a) So,distance traveled by car in "t" seconds:

"u_{car}t=u_{van}t+\frac{1}{2} at^2 ;u_{van}=0"

"15 t=\frac{1}{2}\times a \times t^2"

"a=3 m\/sec ^2"

Solving this equation :

"15 t=\frac{1}{2}\times 3 t^2"

"t=10 \space sec"

(b) Distance traveled by car in "10 \space sec =" "\space u_{car}t=15 \times 10=150 m"